# Rethinking Data Selection at Scale: Random Selection is Almost All You Need

### v1

@@ 1 Introduction
Figure 1: The discrepancy between each methods and random selection on BBH benchmark. The Y-axis represents the differential score, which is computed by subtracting the random selection score from the scores obtained using various methods.
@@
Existing SFT data selection methodologies, both external-scoring and self-scoring, are primarily assessed using several widely recognized IT datasets, such as alpaca-GPT4 (Peng et al., 2023), Dolly (Conover et al., 2023), FLAN (Longpre et al., 2023), WizardLM (Xu et al., 2024), and ShareGPT  (Chiang et al., 2023). These datasets are limited in size and originate from a single source. Meanwhile, in practical applications, to fully leverage the inherent knowledge of LLMs, the SFT process frequently necessitates a substantial volume of varied data, ideally in the range of millions or more. This discrepancy creates a gap between the present SFT data selection strategies and real-world applications. In order to observe the impact brought by the dataset size on the performance of different selection strategies, we analyze the difference in outcomes between existing SFT data selection methods and random selection within source datasets ranging from 10K-30K to 1M on Llama3-8B (AI@Meta, 2024). As shown in Figure 1, when the scale of the datasets increases to 1M, these data selection methods yield suboptimal performance compared with random selection.
@@
Inspired by this finding, we rethink whether SFT data selection methods can work when they are required to handle large-scale IT datasets. For external-scoring approaches, it is impractical to apply them to tackle vast amounts of IT data due to the substantial costs (Liu et al., 2023), we hence focus on the self-scoring methods. For self-scoring approaches, we refer to the article Qin et al. (2024) to categorize the techniques into two types: data quality-based methods and data diversity-based methods. To explore how self-scoring methods influence LLMs’ performance when dealing with large-scale IT data, we evaluate several recent methods on two benchmarks that contain millions of instances. The findings from our experiments reveal three main points:
@@
  • Most self-scoring data selection techniques do not significantly outperform random selection on large-scale datasets. Even though these self-scoring methods can achieve significant gains on small-scale datasets, their effectiveness will be greatly reduced when the data size increases and the data sources become complex.
@@ 3.1 Quality-based Selections
Next, a gradient datastore of projected low-dimensional gradient features was constructed which can be reused for different target tasks. For training datapoints $\bm{z}$, they computed d-dimensional projection of the LoRA gradient ~$\tilde{\nabla}\ell(\bm{z};\bm{\theta}_{i})=\Pi^{\top}\hat{\nabla}\ell(\bm{z};\bm{\theta}_{i})$, where $\Pi^{\top}$ is computed and applied by memory-efficient online implementation of random projections proposed by Park et al. (2023). For validation datapoint $\bm{z}^{\prime}$, they computed $\tilde{\Gamma}(\bm{z}^{\prime},\cdot)=\Pi^{\top}\hat{\Gamma}(\bm{z}^{\prime},\cdot)$.
Finally, LESS computed $\max_{j}\mathrm{Inf}_{\mathrm{Adam}}(\bm{z},\mathcal{D}_{\mathrm{val}}^{(j)})$ for the training set $\bm{z}$ across all sub-validation sets $\mathcal{D}_{val}$. Then it selected the highest score examples to construct $\mathcal{D}_{train}$.
$\mathrm{Inf}_{\mathrm{Adam}}(\bm{z},\mathcal{D}_{\mathrm{val}}^{(j)})=\sum_{i=1}^{N}\bar{\eta}_{i}\frac{\langle\bar{\nabla}\ell(\mathcal{D}_{\mathrm{val}}^{(j)};\bm{\theta}_{i}),\tilde{\Gamma}(\bm{z},\bm{\theta}_{i})\rangle}{\|\bar{\nabla}\ell(\mathcal{D}_{\mathrm{val}}^{(j)};\bm{\theta}_{i})\|\|\tilde{\Gamma}(\bm{z},\bm{\theta}_{i})\|}$ (1)
@@ 3.2 Diversity-based Selections
$\arg\max_{i\in N_{t}}\min_{j\in P_{t}}\Delta\left(\mathbf{n}_{i},\mathbf{p}_{j}\right)$ (8)
At each step, the input parameters to K-Center-Sampling were the model $M_{t}$, the current training pool $P_{t}$, and $\mathcal{D}_{t}$. The selection function K-Center-Sampling then outputs the new data point $N_{t}$, which was added to the training pool for the next iteration $P_{t+1}$.
@@ 5.1 Baseline Methods vs Random
In this section, we reproduce all baseline methods in experiments involving LLaMA3-8B and Qwen2-7B on OpenHermes2.5, the experimental results are presented in Table 1, and results on WildChat are detailed in Table 2. We assess LLaMA3-8B and Qwen2-7B with and without fine-tuning on the entire dataset. All mentioned SFT data selection methods are employed to select 10,000 samples as described in Section 4.3. We randomly run 5 times and all of the results are provided in the tables. Furthermore, 50,000 samples obtained through various methods are also shown in the Appendix Table 5,  6.
As indicated in Table 1 and 2, it is evident that when dealing with extensive and diverse IT datasets, no data selection techniques consistently outperform random sampling by a substantial margin, which implies that the average score exceeds the random score by more than 1%. In most cases, the results of the baseline method are within the range of the results obtained by 5 random runs, and a few methods are even worse than the worst random result, For instance, when evaluating Cross-Entropy on Qwen2-7B using Openhermes2.5, the average result is a mere 54.02, significantly below the lowest score of 57.04 obtained in the 5 random trials.
Table 1: The overall results(%) on a variety of downstream tasks based on Openhermes2.5 dataset. CODE means HumanEval, Random $n$ denotes the $n$th random selection. Except for fine-tuning with the entire Openhermes dataset, the bold numbers indicate the best score of each part, and the underlined numbers indicate the suboptimal score.
@@
Table 2: The overall results(%) on a variety of downstream tasks based on WildChat dataset. CODE means HumanEval, Random $n$ denotes the $n$th random selection. Except for fine-tuning with the entire Openhermes dataset, the bold numbers indicate the best score of each part, and the underlined numbers indicate the suboptimal score.
@@ 5.2 Quality vs Diversity
Tables 1 and 2 demonstrate that the diversity-based selection strategy outperforms the quality-based one. To examine whether prioritizing diversity over data quality improves data selection, we designed a supplementary experiment by incorporating a K-means clustering process on the OpenHermes dataset. Instead of selecting data based solely on method scores, we choose higher-scoring data within each cluster to boost the final training set’s diversity.
Table 4 illustrates that integrating the K-means clustering with quality-based selection methods enhances the effectiveness for most approaches. Notably, Cross Entropy on both Llama3 and Qwen2 models shows improvement over 5% and 3%, respectively, when K-means is used to diversify the data. This suggests that for a large-scale IT dataset, data diversity holds more significance than data quality. This also clarifies why random selection often outperforms most SFT data selection methods, as the random process preserves the dataset’s original distribution and diversity to the greatest possible extent.
@@ 5.3 Baseline Analysis
In contrast, ZIP does not need GPU resources, but the computing process is greedy. It incrementally adds 100 data at a time to the final training subset. For large data scales, it takes approximately 7 days to select 50,000 data. In addition, ZIP serves as a data selection method that operates independently of the model, meaning that the selected data cannot be adaptively tuned on the basis of the model. As illustrated in Tables 1 and 2, the data chosen by ZIP in OpenHermes perform poorly in both Llama3-8B and Qwen2-7B, whereas the data selected in WildChat exhibit the best performance across these models.
@@ 5.4 Which method is the best?
Based on Table 3, it is evident that using token length as the criterion for data selection generally yields optimal results. Specifically, for Llama3, regardless of whether the data source is OpenHermes or WildChat, the results are significantly superior to those achieved by other methods. In addition, the average score on WildChat ($55.51$) surpasses that obtained by fine-tuning with the entire dataset ($54.58$).
Selecting data by token length can stably obtain a relatively high training benefit, reduce the uncertainty caused by randomness, and reduce costs. This approach is particularly beneficial for BASE language models which generally have limited capabilities, as they tend to derive the most significant benefits from training on longer texts.
@@ Appendix A Appendix
In this section, table 5, 6 includes training results of various methodologies with a training dataset comprising 50,000 entries 5, 6.
Table 5: The comprehensive results (%) on various downstream tasks using OpenHermes. Mention that CODE means Humaneval. Algorithmkm means the algorithm has a Kmeans process, and Randomx denotes the xth random selection. The bold numbers indicate the best avg score of each part, and the underlined numbers indicate the suboptimal score.
@@
Table 6: The comprehensive results (%) on various downstream tasks using WildChat. Mention that CODE means Humaneval. Algorithmkm means the algorithm has a Kmeans process, and Randomx denotes the xth random selection. The bold numbers indicate the best avg score of each part, and the underlined numbers indicate the suboptimal score.

### v2

@@ 1 Introduction
Figure 1: The discrepancy between each methods and random selection on BBH benchmark  (Suzgun et al., 2022). The Y-axis represents the differential score, which is computed by subtracting the random selection score from the scores obtained using various methods.
@@
Existing SFT data selection methodologies, both external-scoring and self-scoring, are primarily assessed using several widely recognized IT datasets, such as alpaca-GPT4 (Peng et al., 2023), Dolly (Conover et al., 2023), FLAN (Longpre et al., 2023), WizardLM (Xu et al., 2024), and ShareGPT  (Chiang et al., 2023). These datasets are limited in size and originate from a single source. However, during the SFT stage, a substantially larger scale of data, typically ranging from hundreds of thousands to even millions in size, is frequently necessary. For example, Qwen2 (qwe, 2024) utilized over 500,000 pieces of data during the SFT process. Therefore, in practical applications, in order to fully utilize the inherent knowledge of LLMs, large-scale instruction-following data is essential in the SFT process. Moreover, large-scale data sources not only require a sufficient amount of data, but should also have diverse data sources, such as annotated by professional workers, sourced from real users, or synthesized by models, and rich data types include code data, math data, conversation content, knowledge Q&A, etc.. This discrepancy creates a gap between the present SFT data selection strategies and real-world applications. In order to observe the impact brought by the dataset size on the performance of different selection strategies, we analyze the difference in outcomes between existing SFT data selection methods and random selection within source datasets ranging from 10K-30K to 1M on Llama3-8B (AI@Meta, 2024). As shown in Figure 1, when the scale of the datasets increases to 1M, these data selection methods yield suboptimal performance compared with random selection. Here, “Data size 10K-300K” refers to the data sources used in the original papers of different methods. “Data size 1M” refers to Openhermes2.5-1M dataset (Teknium, 2023).
@@
Inspired by this finding, we rethink whether SFT data selection methods can work when they are required to handle large-scale IT datasets. For external-scoring approaches, it is impractical to apply them to tackle vast amounts of IT data due to the substantial costs (Liu et al., 2023), we hence focus on the self-scoring methods. For self-scoring approaches, we refer to the article Qin et al. (2024) to categorize the techniques into two types: data quality-based methods and data diversity-based methods. Data quality-based methods imply that the approach lays greater emphasis on devising an algorithm and evaluation metrics to compute the score of each data item. Subsequently, the selection is carried out based on the data scores. In contrast, the data diversity-based method is more centered around the diversity of the dataset. To explore how self-scoring methods influence LLMs’ performance when dealing with large-scale IT data, we evaluate several recent methods on two benchmarks that contain millions of instances. The findings from our experiments reveal three main points:
@@
  • Most self-scoring data selection techniques do not significantly outperform random selection on large-scale datasets. Even though these self-scoring methods can achieve significant gains on small-scale datasets, their effectiveness will be greatly reduced when the data size increases and the data sources become complex. While the performance of certain methods does exhibit a marginal edge over the random approach when implemented on particular LLMs, a comprehensive consideration of the trade-off between effectiveness and efficiency leads us to the conclusion that, when dealing with extensive data sources, random selection stands out as the most preferable and advantageous option.
@@ 3.1 Quality-based Selections
Next, a gradient datastore of projected low-dimensional gradient features was constructed which can be reused for different target tasks. For training datapoints $\bm{x}$, they computed d-dimensional projection of the LoRA gradient ~$\tilde{\nabla}\ell(\bm{x};\bm{\theta}_{i})=\Pi^{\top}\hat{\nabla}\ell(\bm{x};\bm{\theta}_{i})$, where $\Pi^{\top}$ is computed and applied by memory-efficient online implementation of random projections proposed by Park et al. (2023). For validation datapoint $\bm{x}^{\prime}$, they computed $\tilde{\Gamma}(\bm{x}^{\prime},\cdot)=\Pi^{\top}\hat{\Gamma}(\bm{x}^{\prime},\cdot)$, where $\tilde{\Gamma}(\bm{x}^{\prime},\cdot)$ represents the gradient values of different data $\bm{x}^{\prime}$ under different optimization states $\cdot$.
Finally, LESS computed $\max_{j}\mathrm{Inf}_{\mathrm{Adam}}(\bm{x},\mathcal{D}_{\mathrm{val}}^{(j)})$ for the training set $\bm{x}$ across all sub-validation sets $\mathcal{D}_{val}$. Then it selected the highest score examples to construct $\mathcal{D}_{train}$.
$\mathrm{Inf}_{\mathrm{Adam}}(\bm{x},\mathcal{D}_{\mathrm{val}}^{(j)})=\sum_{i=1}^{N}\bar{\eta}_{i}\frac{\langle\bar{\nabla}\ell(\mathcal{D}_{\mathrm{val}}^{(j)};\bm{\theta}_{i}),\tilde{\Gamma}(\bm{x},\bm{\theta}_{i})\rangle}{\|\bar{\nabla}\ell(\mathcal{D}_{\mathrm{val}}^{(j)};\bm{\theta}_{i})\|\|\tilde{\Gamma}(\bm{x},\bm{\theta}_{i})\|}$ (1)
@@ 3.2 Diversity-based Selections
$\arg\max_{i\in X_{t}}\min_{j\in P_{t}}\Delta\left(\bm{x_{i}},\bm{p}_{j}\right)$ (8)
At each step, the input parameters to K-Center-Sampling were the model $M_{t}$, the current training pool $P_{t}$, and $\mathcal{D}_{t}$. The selection function K-Center-Sampling then outputs the new data point $X_{t}$, which was added to the training pool for the next iteration $P_{t+1}$.
@@ 5.1 Baseline Methods vs Random
In this section, we reproduce all baseline methods in experiments involving LLaMA3-8B and Qwen2-7B on OpenHermes2.5, the experimental results are presented in Table 1, and results on WildChat are detailed in Table 3. We assess LLaMA3-8B and Qwen2-7B with and without fine-tuning on the entire dataset. All mentioned SFT data selection methods are employed to select 10,000 samples as described in Section 4.3. We randomly run 5 times and all of the results are provided in the tables. Furthermore, 50,000 samples obtained through various methods are also shown in the Appendix Table 6,  7.
Table 1: The overall results (%) on a variety of downstream tasks based on Openhermes2.5 dataset. CODE means HumanEval, Random $n$ denotes the $n$th random selection. Except for fine-tuning with the entire Openhermes dataset, the bold numbers indicate the best score of each part, and the underlined numbers indicate the second highest score.
@@
Table 2: The P-values of the significance tests for each method against the results of five rounds of random selection.
Llama3-8B Qwen2-7B
OpenHermes WildChat OpenHermes WildChat
LESS 0.77 0.45 0.80 0.86
IFD 0.85 0.53 0.85 0.68
SelectIT 0.71 0.79 0.60 0.58
Entropy 0.92 0.46 0.78 0.30
Diverse 0.39 0.58 0.37 0.45
zip 0.55 0.36 0.42 0.31
As indicated in Table 1 and 3, it is evident that when dealing with extensive and diverse IT datasets, no data selection techniques consistently outperform random sampling by a substantial margin, which implies that the average score exceeds the random score by more than 1%. In most cases, the results of the baseline method are within the range of the results obtained by 5 random runs, and a few methods are even worse than the worst random result, For instance, when evaluating Cross-Entropy on Qwen2-7B using Openhermes2.5, the average result is a mere 54.02, significantly below the lowest score of 57.04 obtained in the 5 random trials. Besides, We also conducted the Mann-Whitney U test for each method against the results of 5 rounds of random selection. We adopted the right-tailed test approach, with the testing hypothesis being that the scores of each baseline method on different test tasks are greater than those of the random method. We reported the p-value for each method being significantly better than that of the random method in table 2. We found that the p-values of all methods is higher than 0.05, which indicates that the results of all baseline methods are not greater than those of the random method.
Table 3: The overall results (%) on a variety of downstream tasks based on WildChat dataset. CODE means HumanEval, Random $n$ denotes the $n$th random selection. Except for fine-tuning with the entire Openhermes dataset, the bold numbers indicate the best score of each part, and the underlined numbers indicate the second highest score.
@@ 5.2 Quality vs Diversity
Tables 1 and 3 demonstrate that the diversity-based selection strategy outperforms the quality-based one. To examine whether prioritizing diversity over data quality improves data selection, we designed a supplementary experiment by incorporating a K-means clustering process on the OpenHermes dataset. Instead of selecting data based solely on method scores, we choose higher-scoring data within each cluster to boost the final training set’s diversity.
Table 5 illustrates that integrating the K-means clustering with quality-based selection methods enhances the effectiveness for most approaches. Notably, Cross Entropy on both Llama3 and Qwen2 models shows improvement over 5% and 3%, respectively, when K-means is used to diversify the data. This suggests that for a large-scale IT dataset, data diversity holds more importance than data quality. This also clarifies why random selection often outperforms most SFT data selection methods, as the random process preserves the dataset’s original distribution and diversity to the greatest possible extent.
@@ 5.3 Baseline Analysis
In contrast, ZIP does not need GPU resources, but the computing process is greedy. It incrementally adds 100 data at a time to the final training subset. For large data scales, it takes approximately 7 days to select 50,000 data. In addition, ZIP serves as a data selection method that operates independently of the model, meaning that the selected data cannot be adaptively tuned on the basis of the model. As illustrated in Tables 1 and 3, the data chosen by ZIP in OpenHermes perform poorly in both Llama3-8B and Qwen2-7B, whereas the data selected in WildChat exhibit the best performance across these models.
@@ 5.4 Which method is the best?
Based on Table 4, it is evident that using token length as the criterion for data selection generally yields optimal results. Specifically, for Llama3, regardless of whether the data source is OpenHermes or WildChat, the results are superior to those achieved by other methods. In addition, the average score on WildChat ($55.51$) surpasses that obtained by fine-tuning with the entire dataset ($54.58$). Since random selection may not ensure the best fine-tuning results, we believe that selecting data by token length can stably obtain a relatively high training benefit, reduce the uncertainty caused by randomness, and reduce costs. This approach is particularly beneficial for BASE language models which generally have limited capabilities, as they tend to derive the most significant benefits from training on longer texts.
@@ Appendix A Appendix
In this section, table 6, 7 includes training results of various methodologies with a training dataset comprising 50,000 entries 6, 7.
Table 6: The comprehensive results (%) on various downstream tasks using OpenHermes. Mention that CODE means Humaneval. Algorithmkm means the algorithm has a Kmeans process, and Randomx denotes the xth random selection. The bold numbers indicate the best avg score of each part, and the underlined numbers indicate the second highest score.
@@
Table 7: The comprehensive results (%) on various downstream tasks using WildChat. Mention that CODE means Humaneval. Algorithmkm means the algorithm has a Kmeans process, and Randomx denotes the xth random selection. The bold numbers indicate the best avg score of each part, and the underlined numbers indicate the second highest score.